use the five exponents of exponents simplify

(Y5)3(Y3)2/(Y4)4

Answer by checkley71(8403)

(Y^5)^3=Y^15
(Y^3)^2=Y^6
(Y^4)^4=Y^16
Y^15*Y^6=Y^21
Y^21/Y^16
Y^21*Y^(-16)=Y^5 ANSWER.